OpenStudy (anonymous):

Please help... A plane heads west (180.0 degrees) at a speed of 400.0 mph. The wind's velocity is southeast (315.0 degrees) at 30. mph. What is Ax, the magnitude of the original vector of the plane?

OpenStudy (anonymous):

Let resultant vector R, Wind vector W, A is Plane vector: A = R - W Ax = Rx - Wx = 400 * cos 180 - 30 cos 45 = -400 - 30√2/2 = - 421.21